To solve for -6(-2+n)=102, I know -6x-2=12, but when I solve for -6 times positive n would the equation then be -6n+12 or 12-6n?
Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(52799)

.

-6n +12  is  THE  SAME  as  12-6n.

Visually different, but algebraically equivalent.



Answer by math_tutor2020(3817)


-6n+12 is the same as 12 + (-6n) or simply 12-6n

Think of the template a+b = b+a which says we can add two numbers in any order.
In this case a = -6n and b = 12.


To solve -6(-2+n) = 102, you could distribute to get 12-6n = 102
Or
You could divide both sides by -6 to get -2+n = -17

I'll let the student finish up.